# Gaussian Process Tilted Nonparametric Density Estimation using Fisher Divergence Score Matching

### v1

@@ V Conclusion
We have proposed a GP-tilted density for low dimensional problems. Our Gaussian process makes use of the random Fourier feature representation for tractable nonparametric learning. We propose three Fisher divergence based learning algorithms, for which the GP-RFF score function is mathematically equivalent to a single layer neural network with known hidden layer parameters. Two features and contributions of our methods include:
@@
    We propose a variational approximation to the model posterior using an ELBO-like function based on the Fisher divergence and variational tempering. The corresponding predictive distribution on the data integrates out uncertainty in the model parameter rather than by adding noise to the data, giving an alternative to noise-conditional Fisher divergence methods for our model.
In future work, we plan to adapt our GP-tilted score model and associated algorithms to real world applications and investigate methods for learning the important regularization parameters $\lambda$ (GP prior) and $\eta$ (variational tempering) and kernel width $\gamma$. While we provided suggestions for setting $\eta$ as a function of $\gamma$, these parameter settings were still empirically hand tuned and can be further optimized.

### v2

@@ V Conclusion
We proposed a GP-tilted density for low dimensional estimation problems. Our Gaussian process makes use of the random Fourier feature representation for tractable nonparametric learning. We derived three Fisher divergence based learning algorithms, for which the GP-RFF score function is mathematically equivalent to a single layer neural network with known hidden layer parameters. Two features and contributions of our methods include:
@@
    We propose a variational approximation to the model posterior using an ELBO-like function based on the Fisher divergence and variational tempering. The corresponding predictive distribution on the data integrates out uncertainty in the model parameter rather than by adding noise to the data, giving an alternative to the noise conditional Fisher divergence for our model.
Potential future developments include adapting and applying TGP within larger AI systems for real world applications and investigating methods for learning the important regularization parameters $\lambda$ (GP prior) and $\eta$ (variational tempering), and the kernel width $\gamma$. While we provided suggestions for setting $\eta$ as a function of $\gamma$, parameter settings were still empirically hand tuned and can be further optimized. Other neural network based density estimators may also benefit from incorporating the proposed GP learning framework [zhang2019random, xu2024sparse].